Circular No. 6421

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1996ah IN NGC 5640 J. Mueller reports her discovery of a supernova (mag ∼ 18) located 4′′.7

west and 3′′.3 south of the center of NGC 5640 (α = 14^{h}20^{m}.8, δ = +80^{o}06′, equinox 2000.0). Mueller found SN 1996ah on a IIIa-F plate taken by K. M. Rykoski and herself on June 6 with the 1.2-m Oschin Schmidt telescope in the course of the second Palomar Sky Survey.

P. Garnavich, P. Challis, A. Riess, and R. Kirshner, Harvard-Smith-sonian Center for Astrophysics, report: “A CCD image obtained with the F. L. Whipple Observatory’s 1.2-m telescope on June 17.3 UT confirms the presence of SN 1996ah, 5′′ west and 1′′ south of the galaxy nucleus. A spectrum taken with the Multiple Mirror Telescope on June 19.3 shows that this is a type-Ia supernova between two and three weeks past maximum.”

COMET C/1995 O1 (HALE-BOPP) N. Biver, D. Bockelée-Morvan, P. Colom, J. Crovisier, E. Gérard,

and H. Rauer, Observatoire de Paris-Meudon; D. Despois, Observatoire de Bordeaux; R. Moreno and G. Paubert, Institut de Radio Astronomie Millimétrique (IRAM), Granada; and J. K. Davies and W. R. F. Dent, Joint Astronomy Centre, Hilo, report the first detections of CS and H_{2}CO and observations of CO, HCN, CH_{3}OH and H_{2}S lines in comet C/1995 O1 with the IRAM 30-m telescope performed on June 14–17: “We observed integrated main-beam brightness temperatures of 0.18 ± 0.03 (CO, 115 GHz); 0.69 ± 0.014 (CO, 230 GHz); 0.76 ± 0.02 (HCN, 89 GHz); 1.20 ± 0.03 (H_{2}S, 168 GHz); 0.085 ± 0.011 (H_{2}CO, 225 GHz); and 0.10 ± 0.007 (CS, 147 GHz), in units of K km s^{−1}. The rotational temperatures observed for CH_{3}OH are 14 ± 1 K from lines around 145 GHz and 26 ± 5 K from lines around 157 GHz. The estimated production rates in units of 10^{28} molecules/s are: CO, 8; HCN, 0.07; CH_{3}OH, 0.6; H_{2}S, 0.5; H_{2}CO, 0.010 (assuming parent distribution); and CS, 0.014 (assuming daughter distribution with parent scalelength of 8000 km). At the same period, the OH-production rate from 18-cm observations at the Nançay radio telescope was 12.3 ± 1.3 × 10^{28} s^{−1}.”

Visual m_{1} estimates: June 13.95 UT, 6.3 (K. Hornoch, Lelekovice, Czech Republic, 10×80 binoculars); 15.00, 6.3 (R. J. Bouma, Groningen, The Netherlands, 10×50 binoculars); 16.32, 6.3 (C. S. Morris, Lockwood Valley, CA, 10×50 binoculars); 19.00, 6.2 (R. Haver, Frasso Sabino, Italy, 15×80 binoculars).

1996 June 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green